OpenStudy (anonymous):

Solve using the quadratic formula: 6x2 + 5x = 21

OpenStudy (anonymous):

x = -7/3 or x= 3/2

OpenStudy (anonymous):

subtract 21 and then use the formula. http://upload.wikimedia.org/wikipedia/en/math/8/e/4/8e4fef5352eb498b3534af481c8c4fd4.png